OpenStudy (anonymous):

5/7x+2/7y=1 6x+y=11 place your answer in the form of (x,y)

OpenStudy (mathmale):

5/7x+2/7y=1\[5/7x+2/7y=1\] really should be written as\[\frac{ 5 }{ 7 }x+\frac{ 2 }{ 7 }y=1\] for greater clarity. It'd be much easier if we were to eliminate fractions in this equation. The LCD of this equation is 7. So , please multiply each term by 7 to eliminate fractions.

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

Candy, "OK" is not a proper or helpful response here. I'm asking you to do something; if you can, please do it. Multiply each term by 7 and write out the result\[7(\frac{ 5 }{ 7 }x+\frac{ 2 }{ 7 }y=1)\]

OpenStudy (anonymous):

please help

OpenStudy (mathmale):

I don't understand your situation at all, really. Why is it (apparently) so hard for you to multiply (5/7) by 7?

OpenStudy (anonymous):

yes it is

OpenStudy (mathmale):

Please re-read my question. You didn't answer my question.

OpenStudy (anonymous):

i'm blind

OpenStudy (mathmale):

You can see well enough to respond quickly to whatever I type. If you are willing and able to hold up your end of a conversation meant to help you solve a problem, fine, let's do that. I need your involvement. "OK" simply doesn't cut it as an answer. Please go back and multiply that equation by 7 as I asked you to do earlier.

OpenStudy (mathmale):

Multiply each term within parentheses by 7:\[7(\frac{ 5 }{ 7 }x+\frac{ 2 }{ 7 }y=1)\] Result: 5x + 2y = 7 Your other equation is 6x+y=11 This system of linear equations is slightly easier to solve because the fractions are gone. What must you do next to find a solution?